EDITORIAL

Terry Goodman

This issue of the Missouri Journal of Mathematical Sciences marks the end of our first year of publication. It has been at times, hectic and difficult. Most often, however, it has been challenging, exciting, and rewarding. The response to MJMS has been most positive and we look forward to continued growth.

Special thanks must go to those who have submitted manuscripts to MJMS. Their contributions to the Journal have been significant and we are quite proud of the quality of the articles that have resulted from these manuscripts. We encourage you to contribute in this way to the Journal. We are particularly looking for articles in computer science as well as articles from high school mathematics teachers.

Second, we want to thank those who have served as referees for MJMS. Each has been thorough, prompt, and very professional. We appreciate their efforts. We are still in need of additional referees and we encourage you to contact Robert Kennedy or Bill Grimes. A referee’s application form is found on page 43.

Finally, this must be our last “blanket” mailing. Starting with our next issue in October we will be mailing MJMS only to current subscribers. Therefore, we encourage you who have not ordered a subscription to do so. A subscription order form is found on page 42. We believe that MJMS has had a solid first year and that it has great potential for the future. We hope that you will be a part of this effort to exchange and discuss ideas.

We welcome your comments to us. On page 38 you will find a letter from Andrea Rothbart in response to the editorial in the last issue. She raises some very relevant and interesting points. We encourage this kind of dialogue because we feel that in this way we can all hear more points of view, ideas, and information.

We wish all of you a very enjoyable and productive summer and we look forward to sharing with you again in our October issue of MJMS.